### v2

# One Small Step with Fingerprints, One Giant Leap for De Novo Molecule Generation from Mass Spectra

Abstract
A common approach to the de novo molecular generation problem from mass spectra involves a two-stage pipeline: (1) encoding mass spectra into molecular fingerprints, followed by (2) decoding these fingerprints into molecular structures. In our work, we adopt MIST (Goldman et al., 2023) as the encoder and MolForge (Ucak et al., 2023) as the decoder, leveraging pretraining to enhance performance. Notably, pretraining MolForge proves especially effective, enabling it to serve as a robust fingerprint-to-structure decoder. Additionally, instead of passing the probability of each bit in the fingerprint, thresholding the probabilities as a step function helps focus the decoder on the presence of substructures, improving recovery of accurate molecular structures even when the fingerprints predicted by MIST only moderately resembles the ground truth in terms of Tanimoto similarity. This combination of encoder and decoder results in a tenfold improvement over previous state-of-the-art methods, generating top-1 28% / top-10 36% of molecular structures correctly from mass spectra. We position this pipeline as a strong baseline for future research in de novo molecule elucidation from mass spectra.
1 Introduction
Figure 1: A pipeline of using mass spectrum encoders and fingerprint decoders to solve the de novo molecule generation problem.
Mass spectrometry (MS) is a foundational technique in analytical chemistry, widely employed for the structural characterisation of organic molecules (Noriega et al., 2022). By providing high-resolution measurements of molecular mass (with precision up to four decimal places), MS enables accurate inference of molecular formulas. Additionally, the ionisation process produces fragmentation patterns that encode rich structural information, offering insights into the underlying functional groups and bond connectivity of the molecule. These characteristics make MS an indispensable tool in domains such as metabolomics (Alseekh et al., 2021) and environmental chemistry (Lebedev, 2013), where the identification of previously unseen compounds is critical. Traditional retrieval-based methods, paired with curated spectral databases, have substantially improved in identifying structures from MS data (Bushuiev et al., 2024). However, more needs to be done when dealing with the problem of novel compounds not available in databases. This sets the stage for machine learning approaches to play a central role in de novo molecule generation.
De novo molecule generation from mass spectra has been an elusive problem over the years, but recent work has shown progress towards solving the problem. Progress on machine learning efforts has also been spurred on by curated datasets and benchmarks which has helped to consolidate and compare methods consistently. One of the methods to solve the problem of de novo molecule generation from mass spectra has been to use a two-stage pipeline of (1) encoding mass spectra into molecular fingerprints (FP), followed by (2) decoding these fingerprints into molecular structures (Bohde et al., 2025). An example pipeline is shown in Figure 1. Hence, the problem of de novo molecule generation can be broken down into choosing appropriate encoders and decoders.
Previous approaches utilising this pipeline suffer from poor performance due to the dependency of the fingerprint decoder on the fingerprints generated by the mass spectra encoder. With fingerprints represented as 2048 or 4096 bits, every bit that the encoder is unsure of leads to compounding errors when using the decoder. Additionally, previous decoders may not benefit as much from pretraining due to the architecture of the model.
Our core idea lies in choosing the right fingerprint representation and decoder that mitigates the weaknesses listed above. Since fingerprints represent the presence and absence of substructures in a molecule, we focus on capturing the presence of substructures, since that would be easier for the model to piece together the overall structure. Additionally, a decoder that uses the transformer architecture would be able to scale well with more pretraining data. These two points can be met by using MolForge  (Ucak et al., 2023) as the decoder as our pipeline.
Our key contributions are:
  • We chose MIST (Goldman et al., 2023) for spectrum-to-fingerprint encoding with a pretrained MolForge  (Ucak et al., 2023) decoder for our two-stage pipeline for de novo molecular structure generation from mass spectra.
  • We propose a step-function thresholding of predicted fingerprint probabilities, which focuses representation on the presence of substructures in the molecule. This makes the decoder more robust to noisy or partially correct fingerprints.
  • We demonstrate that pretraining the fingerprint-to-structure decoder significantly enhances generalisation and structure recovery, even when fingerprint similarity to ground truth is moderate.
  • Our method outperforms previous state-of-the-art approaches by an order of magnitude, achieving 28% top-1 and 36% top-10 structure generation accuracy on the MassSpecGym (Bushuiev et al., 2024) dataset.
2 Problem Statement & Definitions
We adopt the problem definition for de novo molecule generation from MassSpecGym (Bushuiev et al., 2024), which involves predicting a molecule’s structure $G$ given its corresponding mass spectrum $MS$. The molecule is modeled as a graph $G=(V,E)$, where the $N$ atoms are represented by the set of vertices $V\in\mathbb{V}^{N}$, and the $M$ chemical bonds are represented by the set of edges $E\in\mathbb{E}^{M}$. The mass spectrum $MS$ consists of intensity values $Y\in(0,1]$ associated with mass-to-charge ratios $X\in\mathbb{R}_{+}$.
This paper focuses on de novo molecule generation with the chemical formula provided as an additional input, which is also part of the MassSpecGym benchmark. In this setting, the chemical formula, which corresponds to the atom set $V$, is known in advance. The task is to predict the molecular structure $G$ given the mass spectrum $MS$ and the chemical formula of the unknown molecule.
Importantly, as this problem is de novo in nature, algorithms attempting to solve this problem need to generate molecular structures without any reference libraries.
2.1 Metrics
Top-$k$ metrics. The de novo molecule generation problem can be formulated as predicting a set of $k$ candidate graphs $\widehat{G}_{k}=\{\widehat{G}_{1},\ldots,\widehat{G}_{k}\}$, rather than a single predicted graph $\widehat{G}$. Especially since mass spectra may not contain enough information to predict structures accurately, this formulation better reflects the uncertainty and complexity of molecule generation from mass spectra.
To evaluate the quality of the predicted molecular graphs $\widehat{G}_{k}$, we compare them against the ground-truth graph $G$ using three metrics, following MassSpecGym (Bushuiev et al., 2024). As a summary:
  1. 1.
    Top-$k$ accuracy.
    $\text{Top-}k\ \text{accuracy:}\ \mathbbm{1}\{G\in\widehat{G}_{k}\}$ (1)
    is the presence of the ground truth molecule within the top-$k$ predictions of the model. This is averaged over all test examples. Here, $\mathbbm{1}\{\cdot\}$ is the indicator function, which returns 1 if the condition is true and 0 otherwise.
  2. 2.
    Maximum Common Edge Subgraph (MCES) metric. (Kretschmer et al., 2025)
[…]
    The MCES metric measures the graph edit distance, or the number of edges that need to be added to candidate structure $\widehat{G}$ such that ground truth $G$ is also a subgraph of it. We report the best similarity score among the top-$k$ candidates as averaged across the test set. A score of 0 indicates identical graphs, while larger values correspond to greater structural dissimilarity.
  3. 3.
    Tanimoto similarity, computed on the Morgan fingerprints of the molecules (Morgan, 1965).
[…]
    This score reflects how well the generated molecule captures true molecular substructures. Tanimoto similarity ranges from 0 to 1, with 1 indicating perfect structural similarity (but not exact similarity, due to the possibility of fingerprint collisions).
2.2 Cheminformatic terms
While exact graph matching can be used to check if the ground truth and predicted molecules match, the cheminformatics domain has also developed other representations for this. In this paper, we focus on using InChIs (Heller et al., 2015) as textual representations of molecules. InChIs are guaranteed to be unique for each structure and thus can be used as a substitute to test for molecule similarity. SMILES strings (Weininger, 1988), though a common text representation of molecules, are not used here, as multiple SMILES strings can be used to refer to the same molecule.
We also use molecular fingerprints as a representation of a molecule, which are usually 2048 or 4096 bits in length. Molecular fingerprints are bit vectors that represent the presence (‘1’ / on-bit) or absence (‘0’ / off-bit) of substructures within a molecule. One of the most commonly used fingerprints is a Morgan fingerprint (Morgan, 1965).
3 Methodology
Our current method utilises a combination of existing models. Here, we show how we have adapted these models to form a pipeline to solve the de novo molecule generation problem. These models (MIST  [^1] https://github.com/samgoldman97/mist/tree/main_v2, MolForge  [^2] https://github.com/knu-lcbc/MolForge) have open source implementations available. We will release source code for our methodology upon acceptance of this paper.
3.1 Mass Spectra Encoder
We use the pretrained MIST models  [^3] https://zenodo.org/records/15122968 from DiffMS. In summary, the MIST model was pretrained on a finetuning dataset to predict molecular fingerprints. The predicted fingerprint is extracted as the output of the spectra encoder in MIST. We refer the reader to Section 3.3 of the DiffMS paper for technical details. It is important to note that during pretraining, MIST does not see any molecular structures that appear in the test set.
3.2 Fingerprint Decoder
MolForge (Ucak et al., 2023) is an autoregressive transformer model that also predicts molecular structure from fingerprints. Specifically, it takes in the on-bits of a fingerprint (bits that have a value of ‘1’) to predict a SMILES string that represents the molecular structure. This is done using an encoder-decoder model that generates a mapping of indices for the encoder and a mapping of tokens (representing atoms, parenthesis, and other symbols in SMILES strings) for the decoder.
MolForge takes the indices of on-bits of the molecular fingerprint as input. However, MIST generates fingerprints as probabilities, real values between 0 to 1. As such, we set a threshold of $t=0.5$ and use all indices of bits $\geq t$ as input to MolForge. This threshold value is a hyperparameter that can be tuned for further optimisation.
MassSpecGym only contains a limited number of molecules in the dataset ($\sim$17k), as there are multiple mass spectra for the same molecule. Thus, there is a need for pretraining such that MolForge has more data to learn the decoding process. We pretrain MolForge with the same dataset used in DiffMS and its baselines, which contains more than 2M compounds and does not include any molecules in the test set, over 6 epochs with a learning rate of $0.0005$.
To generate the top-$k$ molecular predictions, MolForge employs a beam search decoding strategy. At each decoding step, the model expands the most probable partial SMILES sequences by selecting the top-ranked token continuations according to the cumulative log-probability of each sequence. The search proceeds iteratively until complete SMILES strings are formed, after which the top-$k$ highest probability sequences are retained as final outputs.
3.3 Datasets
We are currently using MassSpecGym (Bushuiev et al., 2024) to test our pipeline. For pretraining, we refer to the combined dataset that is made available in DiffMS (Bohde et al., 2025).
There are other datasets, such as the CANOPUS dataset (Dührkop et al., 2021), and we plan to test our pipeline on this as well.
4 Results and Discussion
Table 1: De novo structural elucidation performance on MassSpecGym (Bushuiev et al., 2024) datasets. The best performing model for each metric is bold and the second best is underlined. $\ddagger$} indicates results reproduced from MassSpecGym. * indicates results reproduced from DiffMS.
Top-1 Top-10
Model Accuracy $\uparrow$ MCES $\downarrow$ Tanimoto $\uparrow$ Accuracy $\uparrow$ MCES $\downarrow$ Tanimoto $\uparrow$
MassSpecGym
SMILES Transformer 0.00% 79.39 0.03 0.00% 52.13 0.10
MIST + MSNovelist 0.00% 45.55 0.06 0.00% 30.13 0.15
SELFIES Transformer 0.00% 38.88 0.08 0.00% 26.87 0.13
Spec2Mol 0.00% 37.76 0.12 0.00% 29.40 0.16
MIST + Neuraldecipher 0.00% 33.19 0.14 0.00% 31.89 0.16
Random Generation 0.00% 21.11 0.08 0.00% 18.26 0.11
MADGEN 1.31% 27.47 0.20 1.54% 16.84 0.26
DiffMS 2.30% 18.45 0.28 4.25% 14.73 0.39
Ours (MIST + MolForge) 28.27% 14.72 0.70 36.11% 10.69 0.74
4.1 De novo molecule generation
The primary results of our model pipeline on the de novo molecular generation task are shown in Table 1. We benchmark our pipeline against the state-of-the-art DiffMS model, along with several established baselines reported in prior work. Our approach yields a significant performance gain, achieving an approximately 10-fold increase in exact structure match accuracy when MolForge is used as the fingerprint-to-structure decoder. Furthermore, the Tanimoto similarity between the predicted and ground truth molecular fingerprints exceeds 0.7, indicating strong structural alignment. The Maximum Common Edge Substructure (MCES) metric also shows a reduction in the number of extra atoms required to form a common graph between predicted and reference molecules, implying that predicted structures are more similar to the ground truth structure compared to other methods.
A detailed breakdown of the performance at each stage of our pipeline is provided in the following subsections.
4.2 Performance of MIST as mass spectra encoder
As we have utilised the same pretrained MIST model from the DiffMS paper, we refer the reader to Section 4.4 of the DiffMS paper. DiffMS has also conducted an ablation study, showing that pretraining MIST results in improved performance compared to without pretraining MIST.
MIST also takes in annotations of peaks, in terms of chemical formulae, in the mass spectrum to improve fingerprint prediction. In Section 4.5 of the DiffMS paper, performance of the MIST model is compared between using the true chemical formulae and the chemical formulae provided by MIST-CF. Here, we show in Table 2 that removing annotations from the mass spectra (i.e. not supplying chemical formulae of the peaks into MIST) results in degraded performance.
Table 2: Performance of MIST with/out chemical formulae (CF) annotation of mass spectra peaks. Tanimoto Similarity is calculated between fingerprint of the ground truth molecule and fingerprint predicted by MIST.
MIST model Tanimoto Similarity
With CF 0.731
Without CF 0.627
4.3 Performance of MolForge as fingerprint decoder
Table 3: Performance of MolForge with/out pretraining, as well as the input fingerprint (FP) used. We show that pretraining MolForge leads to vastly improved performance. We also compare ground truth fingerprints, which are generated directly from the ground truth molecular structure, with fingerprints from MIST, which are predicted from the corresponding mass spectra.
Top-1 Top-10
Pretrained MIST? Input FP Accuracy $\uparrow$ Tanimoto $\uparrow$ Accuracy $\uparrow$ Tanimoto $\uparrow$
From MIST 28.27% 0.70 36.11% 0.74
Ground truth 46.00% 0.89 59.28% 0.93
Ground truth 0.00% 0.18 0.00% 0.19
We pretrained MolForge with the same dataset used in DiffMS and its baselines, which contains more than 2M compounds. This allowed it to better capture the distribution of SMILES strings compared to just using the MassSpecGym dataset, which has only $\sim$17k compounds. Without this pretraining, MolForge fails to generate any structures from the test set correctly (see Table 3). The increase in performance with larger pretraining dataset size is expected due to data scaling laws (Kaplan et al., 2020). Hence, we consider this pretraining step to be vital.
It is also important to note that the performance of MolForge with ground truth fingerprints as input is very high with 59% top-10 accuracy (in terms of molecule structural matches) in the test set. This shows that MolForge is a good model for the inverse problem of decoding fingerprints to structure, and why it is a vital component of our pipeline.
Additionally, instead of passing the probability of each bit in the fingerprint, thresholding the probabilities as a step function helps focus the decoder on the presence of substructures. MolForge only takes in the indices of the fingerprint that has a value of 1, representing the presence of probable substructures in the fingerprint. This makes it capable of recovering accurate molecular structures even when the fingerprints predicted by MIST only moderately resembles the ground truth in terms of Tanimoto similarity (0.731, see Table 2).
Hence, the elevated performance of MolForge with ground truth fingerprints shows that the key bottleneck in our pipeline is the prediction of fingerprints from mass spectra. The performance of our pipeline is about half of the best-case scenario if MIST were to predict fingerprints perfectly, suggesting more room for improvement. Even though the fingerprint predicted by MIST has errors, MolForge is able to overcome the noise in the data and generate the correct structure. Further improvements on the prediction of fingerprints from mass spectra should translate well to better performance in the de novo molecule generation problem given the well-rounded performance of MolForge.
5 Relevant Literature
5.1 De novo molecule generation from mass spectra
The problem of de novo molecular generation from mass spectrometry data, also referred to as the inverse problem of structural elucidation from mass spectra, has garnered significant attention in computational chemistry and machine learning. Early approaches leveraged machine learning techniques to map mass spectra to molecular structures (Lim et al., 2018; Wei et al., 2019). However, recent advancements, including the availability of standardised benchmarks and curated datasets, have facilitated more robust evaluation and fairer comparisons across methods.
The MassSpecGym benchmark (Bushuiev et al., 2024) represents one of the most comprehensive open-source datasets for tandem mass spectrometry-based structure elucidation. In addition to the dataset, MassSpecGym introduced a baseline transformer model for the de novo molecule generation problem. Notably, this baseline model failed to correctly predict any structures in the held-out test split, underscoring the intrinsic difficulty of the task.
More recently, DiffMS (Bohde et al., 2025), a generative diffusion model, has been proposed to solve this problem. The method first employs MIST (Goldman et al., 2023) to predict molecular fingerprints from mass spectra. These fingerprints are then used as conditional inputs to a graph diffusion model that generates candidate molecular structures as graphs. This two-stage framework enables DiffMS to achieve a test-set accuracy of 2.30% on the MassSpecGym benchmark, representing a meaningful step forward in this problem.
5.2 Encoding mass spectra to molecular fingerprints
DeepEI (Ji et al., 2020) also encodes a mass spectrum, but does not fully generate a molecular fingerprint. Rather, it identifies salient bits of the fingerprints in their targeted chemical domain, identified as those with a positive percentage between 10% and 90% in the fingerprint training set.
CSI:FingerID (Dührkop et al., 2015) is another model used to generate fingerprints from mass spectra. It uses a fragmentation tree to generate plausible fragments. However, this fingerprint type is different from Morgan fingerprints, and thus is not as extensible for usage with other fingerprint decoder models.
5.3 Decoding molecular fingerprints to molecular structure
DiffMS (Bohde et al., 2025), as stated earlier, contains a conditional graph diffusion model that takes in a fingerprint representation to predict molecule structure. Together with information about the chemical formula, in which the set of atoms are already known, DiffMS is used to predict the connectivity graph, or how atoms are connected to form the molecular structure.
DiffMS has also explored other fingerprint-to-structure models to be paired with MIST. Neuraldecipher (Le et al., 2020) uses a standard feed-forward model to decode the molecular fingerprint, while MSNovelist (Stravs et al., 2022) uses a recurrent neural network with long short-term memory architecture.
6 Conclusion
Various two-stage pipelines have been proposed for solving the de novo molecular generation problem from mass spectra, comprising (1) encoding mass spectra into molecular fingerprints, followed by (2) decoding these fingerprints into molecular structures. Leveraging pretraining on external chemical datasets significantly enhances performance in both stages, especially so for the decoder. This stems from the limited availability of labeled mass spectra, in contrast to the abundance of molecular structure data that can be exploited during decoder training. Notably, substituting the decoder with MolForge, a seemingly modest architectural change, yields a tenfold improvement over prior state-of-the-art, highlighting the how choosing the right decoder is critical. We position this pipeline as a strong baseline for future work in de novo structure prediction from mass spectra, and highlight the spectra-to-fingerprint encoding step as a promising direction for further investigation.