### v3

# One Small Step with Fingerprints, One Giant Leap for De Novo Molecule Generation from Mass Spectra

Abstract
A common approach to the de novo molecular generation problem from mass spectra involves a two-stage pipeline: (1) encoding mass spectra into molecular fingerprints, followed by (2) decoding these fingerprints into molecular structures. In our work, we adopt MIST Goldman u. a. (2023) as the encoder and MolForge Ucak u. a. (2023) as the decoder, leveraging additional training data to enhance performance. We also threshold the probabilities of each fingerprint bit to focus on the presence of substructures. This results in a tenfold improvement over previous state-of-the-art methods, generating top-1 28% / top-10 36% of molecular structures correctly from mass spectra in MassSpecGym Bushuiev u. a. (2024). We position this as a strong baseline for future research in de novo molecule elucidation from mass spectra.
DSO National Laboratories
Singapore
Figure 1: A pipeline of using mass spectrum encoders and fingerprint decoders to solve the de novo molecule generation problem.
1 Introduction
De novo molecule generation from mass spectra has been an elusive problem over the years. One of the methods to solve this has been to use a two-stage pipeline of (1) encoding mass spectra into molecular fingerprints (FP), followed by (2) decoding these fingerprints into molecular structures (Bohde u. a., 2025), shown in Figure 1. Hence, choosing appropriate encoders and decoders is important.
Previous approaches utilising this pipeline suffer from poor performance due to the dependency of the fingerprint decoder on the fingerprints generated by the mass spectra encoder. Every bit that the encoder is unsure of leads to compounding errors when using the decoder.
Our core idea lies in choosing the right decoder that mitigates the weaknesses above. Since fingerprints represent the presence and absence of substructures in a molecule, we focus on capturing only present substructures, making it easier for the model to piece together the overall structure. Additionally, a decoder that uses the transformer architecture would be able to scale well with more training data. These two points can be met by using MolForge  (Ucak u. a., 2023) as the decoder in our pipeline.
Our key contributions are:
  • We chose MIST (Goldman u. a., 2023) for spectrum-to-fingerprint encoding with MolForge  (Ucak u. a., 2023) decoder for our two-stage pipeline for de novo molecular structure generation from mass spectra.
  • We demonstrate that training the fingerprint-to-structure decoder on a larger dataset with varying chemical domains significantly enhances generalisation and structure recovery, even when fingerprint similarity to ground truth is moderate.
  • Our method outperforms previous state-of-the-art by an order of magnitude, achieving 28% top-1 and 36% top-10 structure generation accuracy on the MassSpecGym (Bushuiev u. a., 2024) dataset.
2 Problem Statement & Definitions
We adopt the problem definition and metrics (top-$k$ accuracy, Maximum Common Edge Subgraph (MCES) and Tanimoto similarity) for de novo molecule generation from MassSpecGym (Bushuiev u. a., 2024). Further details can be found in Appendix A. Importantly, as this problem is de novo in nature, algorithms attempting to solve this problem need to generate molecular structures without any reference libraries.
3 Methodology
Our current method utilises a combination of existing models. Here, we show how we have adapted these models to form a pipeline to solve the de novo molecule generation problem. These models (MIST , MolForge ) have open source implementations available. Implementation details can be found in Appendix B.
3.1 Mass Spectra Encoder
We use the pretrained MIST models from DiffMS. The MIST model was trained on MassSpecGym to predict molecular fingerprints from mass spectra, where each peak is annotated with possible chemical formulae by an automated annotation program. We refer the reader to Section 3.3 of the DiffMS paper for technical details. In short, each peak in the mass spectra is encoded via a chemical formula transformer, then passed to a module that outputs the predicted fingerprint.
3.2 Fingerprint Decoder
MolForge (Ucak u. a., 2023) is an autoregressive transformer model that also predicts molecular structure from fingerprints. Specifically, it takes in the on-bits of a fingerprint (bits that have a value of ‘1’) to predict a SMILES string that represents the molecular structure. This is done using an encoder-decoder model that generates a mapping of indices for the encoder and a mapping of tokens (representing atoms, parenthesis, and other symbols in SMILES strings) for the decoder.
MolForge takes the indices of on-bits of the molecular fingerprint as input. However, MIST generates fingerprints as probabilities, real values between 0 to 1. As such, we set a threshold of $t=0.5$ and use all indices of bits $\geq t$ as input to MolForge. This threshold value is a hyperparameter that can be tuned for further optimisation.
MassSpecGym only contains a limited number of molecules in the dataset ($\sim$17k), as there are multiple mass spectra for the same molecule. Thus, there is a need for further training such that MolForge has more data to learn the decoding process. We train MolForge with the same dataset used in DiffMS and its baselines, which contains $\sim$2M compounds and does not include any molecules in the test set, over 6 epochs with a learning rate of $0.0005$. MolForge was trained to take in 4096-bit fingerprints as the pretrained MIST models generate 4096-bit fingerprints.
To generate the top-$k$ molecular predictions, MolForge employs a beam search decoding strategy. At each decoding step, the model expands the most probable partial SMILES sequences by selecting the top-ranked token continuations according to the cumulative log-probability of each sequence. The search proceeds iteratively until complete SMILES strings are formed, after which the top-$k$ highest probability sequences are retained as final outputs.
3.3 Datasets
We use MassSpecGym (Bushuiev u. a., 2024) to test our pipeline. For training on additional compounds, we refer to the combined dataset that is made available in DiffMS (Bohde u. a., 2025), which corresponds to data from DSSTox (CCTE, 2024), HMDB (Wishart u. a., 2022), COCONUT (Sorokina u. a., 2021), and MOSES (Polykovskiy u. a., 2020) datasets. This results in about 2 million unique compounds, and we use this dataset for as additional training for MolForge.
4 Results and Discussion
Table 1: De novo structural elucidation performance on MassSpecGym (Bushuiev u. a., 2024) datasets, across accuracy (Acc), MCES edit distance (MCES) and Tanimoto similarity of fingerprints (Tanimoto).
Top-1 Top-10
Model Acc. $\uparrow$ MCES $\downarrow$ Tanimoto $\uparrow$ Acc. $\uparrow$ MCES $\downarrow$ Tanimoto $\uparrow$
MassSpecGym
MADGEN 1.31% 27.47 0.20 1.54% 16.84 0.26
DiffMS 2.30% 18.45 0.28 4.25% 14.73 0.39
Ours (MIST+MolForge) 28.27% 14.72 0.70 36.11% 10.69 0.74
4.1 De novo molecule generation
The primary results of our model pipeline on the de novo molecular generation task are shown in Table 1. We benchmark our pipeline against the state-of-the-art DiffMS model, along with several established baselines reported in prior work. Our approach yields a significant performance gain, achieving an approximately 10-fold increase in exact structure match accuracy when MolForge is used as the fingerprint-to-structure decoder. Furthermore, the Tanimoto similarity between the predicted and ground truth molecular fingerprints exceeds 0.7. The Maximum Common Edge Substructure (MCES) metric also shows a reduction in the number of extra atoms required to form a common graph between predicted and reference molecules, implying that predicted structures are more similar to the ground truth structure compared to other methods.
4.2 Performance of MIST as mass spectra encoder
As we have utilised the same pretrained MIST model from the DiffMS paper, we refer the reader to Section 4.4 of the DiffMS paper. DiffMS has also conducted an ablation study, showing that pretraining MIST results in improved performance compared to without pretraining MIST.
MIST also takes in annotations, in terms of chemical formulae, of peaks in the mass spectrum. In our study, removing annotations from the mass spectra (i.e. not supplying chemical formulae of the peaks into MIST) results in a decrease in Tanimoto Similarity (as calculated between fingerprint of the ground truth molecule and fingerprint predicted by MIST) from 0.731 to 0.627. Hence, peak annotations of the mass spectra are also essential for the good performance of our proposed pipeline.
4.3 Performance of MolForge as fingerprint decoder
Table 2: Performance of MolForge on different training datasets and the input fingerprint (FP) used. We show that additional training data leads to vastly improved performance for MolForge. We also compare ground truth fingerprints, generated directly from the ground truth structure, with fingerprints from MIST, which are predicted from the corresponding mass spectra.
Top-1 Top-10
MolForge training dataset Input FP Acc. $\uparrow$ Tanimoto $\uparrow$ Acc. $\uparrow$ Tanimoto $\uparrow$
Combined DiffMS ($\sim$2M) From MIST 28.27% 0.70 36.11% 0.74
Combined DiffMS ($\sim$2M) Ground truth 46.00% 0.89 59.28% 0.93
MassSpecGym ($\sim$17k) Ground truth 0.00% 0.18 0.00% 0.19
We trained MolForge with the same dataset used in DiffMS and its baselines, which contains more than two million compounds. This allowed it to better capture the distribution of SMILES strings compared to just using the MassSpecGym dataset, which has only $\sim$17k compounds. Without these additional data, MolForge fails to generate any structures from the test set correctly (see Table 2). The increase in performance with larger dataset size is expected due to data scaling laws (Kaplan u. a., 2020). Hence, we consider this step to be vital.
It is also important to note that the performance of MolForge with ground truth fingerprints as input is very high with 59% top-10 accuracy (in terms of molecule structural matches) in the test set. This shows that MolForge is a good model for the inverse problem of decoding fingerprints to structure, and why it is a vital component of our pipeline.
Additionally, instead of passing the probability of each bit in the fingerprint, thresholding the probabilities as a step function helps focus the decoder on the presence of substructures. MolForge only takes in the indices of the fingerprint that has a value of 1, representing the presence of probable substructures in the fingerprint. This makes it capable of recovering accurate molecular structures even when the fingerprints predicted by MIST only moderately resembles the ground truth in terms of Tanimoto similarity.
Hence, the elevated performance of MolForge with ground truth fingerprints shows that the key bottleneck in our pipeline is the prediction of fingerprints from mass spectra. The performance of our pipeline is about half of the best-case scenario if MIST were to predict fingerprints perfectly, suggesting room for improvement. Even though the fingerprint predicted by MIST has errors, MolForge is able to overcome them and generate the correct structure. Further improvements on the prediction of fingerprints from mass spectra should translate well to better performance in the de novo molecule generation problem given the well-rounded performance of MolForge.
5 Relevant Literature
De novo molecule generation from mass spectra. The problem of de novo molecular generation from mass spectrometry data, also referred to as the inverse problem of structural elucidation from mass spectra, has garnered significant attention in computational chemistry and machine learning. Early approaches leveraged machine learning techniques to map mass spectra to molecular structures (Lim u. a., 2018; Wei u. a., 2019). The MassSpecGym benchmark (Bushuiev u. a., 2024) represents one of the most comprehensive open-source datasets for tandem mass spectrometry-based structure elucidation to compare across models. More recently, DiffMS (Bohde u. a., 2025), a generative diffusion model, has been proposed to solve this problem. The method first employs MIST (Goldman u. a., 2023) to predict molecular fingerprints from mass spectra. These fingerprints are then used as conditional inputs to a graph diffusion model that generates candidate molecular structures as graphs. This two-stage framework enables DiffMS to achieve a test-set accuracy of 2.30% on the MassSpecGym benchmark.
Encoding mass spectra to molecular fingerprints. DeepEI (Ji u. a., 2020) also encodes a mass spectrum, but does not fully generate a molecular fingerprint. CSI:FingerID (Dührkop u. a., 2015) uses a fragmentation tree to generate plausible fragments. However, this fingerprint type is different from Morgan fingerprints, and thus is not as extensible for usage with other fingerprint decoder models.
Decoding molecular fingerprints to molecular structure. DiffMS (Bohde u. a., 2025), as stated earlier, contains a conditional graph diffusion model that takes in a fingerprint representation to predict molecule structure. Other fingerprint-to-structure models can be paired with MIST. Neuraldecipher (Le u. a., 2020) uses a standard feed-forward model to decode the molecular fingerprint, while MSNovelist (Stravs u. a., 2022) uses a recurrent neural network with long short-term memory architecture.
6 Conclusion
Various two-stage pipelines have been proposed for solving the de novo molecular generation problem from mass spectra, comprising (1) encoding mass spectra into molecular fingerprints, followed by (2) decoding these fingerprints into molecular structures. Leveraging external chemical datasets significantly enhances performance in both stages, especially so for the decoder. This stems from the limited availability of labeled mass spectra, in contrast to the abundance of molecular fingerprint data that can be exploited during decoder training. Notably, substituting the decoder with MolForge, a seemingly modest architectural change, yields a tenfold improvement over prior state-of-the-art, highlighting the how choosing the right decoder is critical. We position this pipeline as a strong baseline for future work in de novo structure prediction from mass spectra, and highlight the spectra-to-fingerprint encoding step as a promising direction for further investigation.
Appendix A Problem Statement and Definitions
We adopt the problem definition for de novo molecule generation from MassSpecGym (Bushuiev u. a., 2024), which involves predicting a molecule’s structure $G$ given its corresponding mass spectrum $MS$. The molecule is modeled as a graph $G=(V,E)$, where the $N$ atoms are represented by the set of vertices $V\in\mathbb{V}^{N}$, and the $M$ chemical bonds are represented by the set of edges $E\in\mathbb{E}^{M}$. The mass spectrum $MS$ consists of intensity values $Y\in(0,1]$ associated with mass-to-charge ratios $X\in\mathbb{R}_{+}$.
This paper focuses on de novo molecule generation with the chemical formula provided as an additional input, which is also part of the MassSpecGym benchmark. In this setting, the chemical formula, which corresponds to the atom set $V$, is known in advance. The task is to predict the molecular structure $G$ given its mass spectrum $MS$ and chemical formula.
Importantly, as this problem is de novo in nature, algorithms attempting to solve this problem need to generate molecular structures without any reference libraries.
A.1 Metrics
Top-$k$ metrics. The de novo molecule generation problem can be formulated as predicting a set of $k$ candidate graphs $\widehat{G}_{k}=\{\widehat{G}_{1},\ldots,\widehat{G}_{k}\}$, rather than a single predicted graph $\widehat{G}$. Especially since mass spectra may not contain enough information to predict structures accurately, this formulation better reflects the uncertainty and complexity of molecule generation from mass spectra.
To evaluate the quality of the predicted molecular graphs $\widehat{G}_{k}$, we compare them against the ground-truth graph $G$ using three metrics, following MassSpecGym (Bushuiev u. a., 2024). As a summary:
  1. 1.
    Top-$k$ accuracy.
    $\text{Top-}k\ \text{accuracy:}\ \mathbbm{1}\{G\in\widehat{G}_{k}\}$ (1)
    is the presence of the ground truth molecule within the top-$k$ predictions of the model. This is averaged over all test examples. Here, $\mathbbm{1}\{\cdot\}$ is the indicator function, which returns 1 if the condition is true and 0 otherwise.
  2. 2.
    Maximum Common Edge Subgraph (MCES) metric. (Kretschmer u. a., 2025)
[…]
    The MCES metric measures the graph edit distance, or the number of edges that need to be added to candidate structure $\widehat{G}$ such that ground truth $G$ is also a subgraph of it. We report the best similarity score among the top-$k$ candidates as averaged across the test set. A score of 0 indicates identical graphs, while larger values correspond to greater structural dissimilarity.
  3. 3.
    Tanimoto similarity, computed on the Morgan fingerprints of the molecules (Morgan, 1965).
[…]
    This score reflects how well the generated molecule captures true molecular substructures. Tanimoto similarity ranges from 0 to 1, with 1 indicating perfect structural similarity (but not exact similarity, due to the possibility of fingerprint collisions).
A.2 Cheminformatics terms
While exact graph matching can be used to check if the ground truth and predicted molecules match, the cheminformatics domain has also developed other representations for this. In this paper, we focus on using InChIs (Heller u. a., 2015) as textual representations of molecules. InChIs are guaranteed to be unique for each structure and thus can be used as a substitute to test for molecule similarity. SMILES strings (Weininger, 1988), though a common text representation of molecules, are not used here, as multiple SMILES strings can be used to refer to the same molecule.
We also use molecular fingerprints as a representation of a molecule, which are usually 2048 or 4096 bits in length. Molecular fingerprints are bit vectors that represent the presence (‘1’ / on-bit) or absence (‘0’ / off-bit) of substructures within a molecule. One of the most commonly used fingerprints is a Morgan fingerprint (Morgan, 1965).
Appendix B Implementation Details
Open source implementations are available for the models (MIST  [^1] https://github.com/samgoldman97/mist/tree/main_v2, MolForge  [^2] https://github.com/knu-lcbc/MolForge) in our pipeline. We have generally followed the default settings for these models, and further modifications are explained in this section.
MIST: We use pretrained MIST models from DiffMS  [^3] https://zenodo.org/records/15122968, and base our implementations from the DiffMS code and annotated MassSpecGym data. We proceed with the inference run with the MIST models pretrained for MassSpecGym, and collect the output of the encoder (MIST model) in test_step as the fingerprint outputs. We set a threshold of 0.5 and extracted all indices for which its corresponding bit is at least the threshold value.
MolForge: The training data was formatted following the examples in the MolForge Github repository, in which SMILES strings were split into tokens (i.e. atoms, brackets or numbers) and fingerprints were generated using the Morgan fingerprint generator from RDKit  [^4] https://www.rdkit.org with a radius of 2 and fingerprint size of 4096 bits. We first train the tokenizer for this model with the additional training data as provided in DiffMS. Following which, we trained the model on the same dataset with a learning rate of 5e-4 and batch size of 128 for 6 epochs. This took about 3 days to run with a Nvidia A40 GPU. For inference, we used the beam search code from MolForge with a beam size of 10.